OpenStudy (anonymous):

the principal P is borrowed and the loans future value A at a time t is given. determine the loans simple interest rate r to the nearest tenth of a percent. P=$3300.00 A=$3345.38 T=3months

OpenStudy (anonymous):

@Jhannybean

OpenStudy (anonymous):

@DanJS

OpenStudy (jhannybean):

Simple interest rate: \(A = P(1+rt)\) P = 3300.00 A = 3345.38 t = 3 Now reformat your problem so you're solving for \(r\)

OpenStudy (anonymous):

3345.38=3300.00(1+r 3)

OpenStudy (jhannybean):

Thanks @jim_thompson5910 correction: t = 3 months / 12 months = # of years

OpenStudy (anonymous):

I need help

OpenStudy (danjs):

this one?

OpenStudy (anonymous):

yes

OpenStudy (danjs):

Basic Idea is for a loan in general Future Value = Current Value + Interest

OpenStudy (danjs):

Simple Interest is just the constant rate of increase per time and is Simple Interest = Current Value + Interest Rate * time in years

OpenStudy (danjs):

Using the variable names they give you , you get A = P + P*r*t you can just factor the P A = P(1+r*t)

OpenStudy (danjs):

That is the idea of where that formula comes from, in the future, the loan will be valued as the current value plus interest... this is one of different interests

OpenStudy (danjs):

A = P*(1+r*t) Putting in the numbers 3345.38 = 3300*[ 1 + r*(3/12)]

OpenStudy (danjs):

there is your rate you need, solve

OpenStudy (anonymous):

3345.38=3300+2475

OpenStudy (anonymous):

3345.38=5775

OpenStudy (danjs):

where did the rate go

OpenStudy (anonymous):

that was .25*3

OpenStudy (danjs):

you are solving for the interest rate , r 3345.38 = 3300*[ 1 + r*(3/12)]

OpenStudy (anonymous):

4125*r

OpenStudy (anonymous):

4.055

OpenStudy (danjs):

looks good

OpenStudy (danjs):

that is a number, percents are , a number out of 100

OpenStudy (anonymous):

I'm confused

OpenStudy (danjs):

3345.38 = 3300*[ 1 + r*(3/12)] 3345.38 = 3300 + 3300 * 0.25 * r 45.38 = 3300*0.25 * r 45.38 = 825*r r = 45.38/825 or about 0.055

OpenStudy (danjs):

sorry i typed it wrong last time , tha tis right

OpenStudy (anonymous):

it says to round to the nearest tenth of percent

OpenStudy (danjs):

that value for r is a decimal, you have to make it a percent

OpenStudy (anonymous):

5%?

OpenStudy (danjs):

yeah, to convert a decimal number to a percent, just multiply by 100, move decimal right 2 places... r = 5.5%

OpenStudy (anonymous):

thank you

OpenStudy (danjs):

welcome